# TenonOS: A Self-Generating Intelligent Embedded Operating System Framework for Edge Computing

### v1

@@ Abstract
The rapid evolution of edge computing has exposed fundamental limitations in traditional operating system and hypervisor architectures, particularly in managing heterogeneous platforms and meeting the constraints of limited resources. Existing solutions often rely on monolithic or layered combinations of hypervisors and guest OSes, which are difficult to tailor for the diverse and dynamic requirements of edge scenarios.
To address these challenges, we propose TenonOS, a demand-driven, self-generating, and lightweight operating system framework that fundamentally rethinks and reconstructs both the hypervisor and OS architectures. TenonOS introduces a novel LibOS-on-LibOS approach, in which both virtualization and OS functionalities are modularized into fine-grained, reusable micro-libraries. A dynamic orchestration engine composes these modules on demand to construct customized, application-specific runtime environments. At the core of TenonOS are two key components: Mortise, a minimal, modularized hypervisor, and Tenon, a real-time LibOS. Mortise provides low-overhead resource isolation, fast inter-VM communication, and manages the full lifecycle of Tenon instances—including on-demand creation, suspension, and termination—enabling TenonOS to flexibly adapt its runtime layout to workload variations. Tenon delivers deterministic scheduling and multi-process support for time-critical applications. Through this unified and modular architecture, TenonOS eliminates redundant layers, reduces system overhead, and enhances scalability, security, and maintainability. Extensive evaluations demonstrate that TenonOS achieves superior real-time scheduling (40.28% improvement), a compact memory footprint (361 KiB), and high adaptability to dynamic edge workloads, making it an ideal foundation for heterogeneous, resource-constrained edge systems.
@@
Edge computing, LibOS, hypervisor, real-time OS
@@ 1 Introduction
Operating systems (OS) serve as the backbone of the information industry, acting as the essential interface between hardware and software. At the foundational level, OSs manage and coordinate hardware resources, while at a higher level, they enable the development and flourishing of complex application ecosystems [1, 2]. The enduring success of operating systems relies on their ability to create robust, adaptable, and scalable environments for both developers and users. However, the accelerating pace of technological innovation and the explosive growth of interconnected devices are pushing the limits of traditional OS architectures [3, 4].
@@
These limitations become especially pronounced in the context of edge computing, where environments are characterized by high dynamism, heterogeneity, and resource constraints [5, 6]. The proliferation of smart devices, sensors, and interconnected systems has led to a fragmented and diverse network landscape [7]. In edge scenarios, computation, storage, and data processing are performed closer to the data source to satisfy requirements for low latency, high bandwidth, and real-time responsiveness [6]. Edge computing thus underpins a wide range of critical applications, including autonomous vehicles, industrial automation, and smart cities. Operating systems in these environments must manage heterogeneous hardware platforms, guarantee strong security and isolation, and operate efficiently under stringent resource limitations—requirements that far exceed those of traditional, centralized systems. Consequently, conventional OS architectures, originally designed for stable hardware and centralized computing, often fail to adapt to the dynamic and fragmented nature of the edge [8, 9].
@@
A particularly pressing challenge in edge environments is the need for multi-OS collaboration. Many edge applications require multiple operating systems to coexist and cooperate on a single device. For example, autonomous vehicles commonly rely on a real-time operating system (RTOS) to handle safety-critical functions, while simultaneously running a general-purpose operating system (GPOS) for infotainment and user applications [10, 11, 12, 13, 14]. This trend highlights the growing demand for efficient multi-OS management and coordination in increasingly diverse and complex edge settings.
@@
To address these needs, virtualization technologies have become a mainstream solution, enabling multiple isolated environments to run on shared hardware. By abstracting the underlying physical resources, virtualization improves flexibility and scalability, and provides strong isolation between different OS instances [6, 15]. However, virtualization is not without its drawbacks. Different layers—such as hypervisors, container runtimes, and guest operating systems—often employ independent resource management strategies, leading to inefficiencies. These inconsistencies can cause resource contention, increased latency, and degraded overall system performance, which are particularly problematic for latency-sensitive and resource-constrained edge applications [16, 17, 18, 19, 20, 21]. Moreover, the overhead and complexity introduced by hypervisor-based solutions may compromise real-time guarantees and make system design more challenging, especially when supporting a wide variety of hardware configurations.
@@
To tackle these challenges, we propose an innovative solution: TenonOS, a self-generating, intelligent, and lightweight operating system framework for edge computing. TenonOS is built on a novel LibOS-on-LibOS architecture that rethinks the conventional separation between the operating system and the hypervisor. Adopting a unified library-first design, TenonOS decomposes both layers into modular micro-libraries, with every functionality implemented as a minimal, composable library. By supporting dynamic composition at runtime, TenonOS provides a flexible and efficient execution environment. This approach eliminates much of the redundancy found in traditional OS-hypervisor stacks, enabling more compact, scalable, and hardware-aware deployments.
@@
The core of TenonOS is its two-layer architecture, composed of the hypervisor Mortise and the runtime Tenon. This structure draws inspiration from the natural growth of plants: Mortise serves as the fertile soil of the system, while Tenon acts as a seed from which multiple distinct operating systems can be generated.
@@
Mortise is a lightweight hypervisor that separates traditional virtualization functions into modular micro-libraries. It supports efficient hardware resource management, low-latency inter-VM communication, and dynamic allocation policies. Tenon, built atop Mortise, is a real-time capable LibOS designed to meet the timing constraints of modern embedded applications. In contrast to conventional LibOS designs, Tenon integrates real-time scheduling mechanisms that support both cooperative and preemptive models, enabling deterministic task execution for critical workloads. It also supports multi-process execution, allowing compatibility with general-purpose applications while maintaining a compact, modular structure. The Mortise hypervisor can concurrently host both Tenon and general-purpose operating systems, such as Linux, on the same hardware platform. This architecture supports mixed-criticality workloads and reduces the need for isolated compute nodes.
@@
A key feature of Mortise is its ability to manage the full lifecycle of Tenon instances. This includes on-demand instantiation, suspension, and termination of LibOS-based runtimes based on system workload or application requirements. By enabling fine-grained control over when and how operating system instances are created and destroyed, Mortise provides the foundation for TenonOS’s self-generating capability. Rather than relying on static runtime environments, TenonOS can dynamically assemble, launch, and retire execution contexts as needed—allowing the system to elastically adapt to workload variations while minimizing resource consumption.
@@
Together, Mortise and Tenon form a cohesive framework that bridges hypervisor-level resource control with application-level runtime flexibility. Mortise handles low-level isolation and scheduling, while Tenon delivers customizable execution environments optimized for responsiveness and scalability. This design aligns well with the diverse and performance-sensitive demands of edge computing deployments.
This paper presents the design and implementation of TenonOS, with the following key contributions:
  • A unified LibOS-based architecture that fundamentally rethinks and reconstructs both operating system and hypervisor frameworks, enabling seamless integration across layers.
  • A modular micro-library design that promotes code reuse, enhances maintainability, and strengthens system security.
  • A unified resource management strategy that eliminates redundant layers and optimizes overall system performance and resource utilization.
  • A high-efficiency, real-time capable LibOS that delivers deterministic scheduling and low-latency operations, thus meeting the stringent performance requirements of time-sensitive applications.
  • Full source code released as open source at https://gitee.com/tenonos/tenon.git and https://gitee.com/tenonos/mortise.git, supporting reproducibility and community collaboration.
2 Background
@@
Edge computing has emerged as a critical paradigm in modern computing infrastructure, driven by the exponential growth of edger devices and the increasing demand for low-latency applications. This distributed computing model brings computation and data storage closer to the data source, enabling real-time processing for applications such as autonomous vehicles, industrial automation, and smart cities. The unique characteristics of edge environments - including heterogeneous hardware architectures, resource-constrained devices, and distributed deployments - present significant challenges for traditional computing approaches.
@@
The evolution of operating systems has historically been shaped by the needs of centralized computing environments. Traditional operating system architectures, developed during the mainframe and personal computing eras, were designed with assumptions of relatively homogeneous hardware and abundant resources. These systems employ monolithic or modular designs that abstract hardware resources through well-defined interfaces, providing process management, memory allocation, and device control functionalities. While effective in conventional computing environments, these architectures face limitations when applied to edge computing scenarios.
Virtualization technology, originally developed for server consolidation in data centers, has been adapted for edge computing environments. Type-1 hypervisors, which run directly on hardware, provide strong isolation between virtual machines and enable efficient resource partitioning. However, the traditional virtualization stack introduces performance overhead due to multiple abstraction layers and redundant resource management between the hypervisor and guest operating systems. This overhead becomes particularly problematic in resource-constrained edge devices where computational resources are limited and latency requirements are strict.
3 Motivation
Modern edge computing applications demand low-latency, real-time and high-throughput responsiveness [22, 23]. To meet these demands, virtualization techniques such as VMs and containers are often adopted for workload isolation and resource sharing [15]. However, these solutions struggle to deliver efficient performance in heterogeneous, resource-constrained edge environments [24, 25].
@@ 3.1 Issues with Monolithic Architectures
In traditional monolithic operating system and virtualization architectures, all core functionalities are tightly coupled within a large, complex codebase. While this design was sufficient in earlier eras with relatively homogeneous and stable hardware environments, it exposes significant drawbacks in today’s heterogeneous and rapidly evolving edge scenarios. First, the **development and maintenance costs are high**, as the tightly intertwined modules make updates or bug fixes require large-scale modifications across the system, thereby increasing the risk of introducing new vulnerabilities. Second, monolithic systems are difficult to **tailor or optimize for specific applications or hardware platforms**, resulting in inefficient resource utilization—an issue that is particularly critical for resource-constrained edge devices. Finally, the monolithic design lacks the flexibility needed to **adapt to diverse accelerators and dynamic workload requirements**, limiting scalability and responsiveness. Consequently, monolithic architectures are increasingly inadequate for addressing the heterogeneity, dynamism, and real-time demands of modern edge computing environments.
@@ 3.2 The Case for Virtualization-OS Structures
A key limitation stems from the layered architecture of traditional systems, where both the hypervisor and the OS independently manage hardware resources. This separation leads to duplicated functionality and conflicting policies. For example, virtual memory is maintained both at the hypervisor and guest OS levels [26], and CPU scheduling occurs twice—first by the hypervisor for vCPUs, then by the OS for processes—resulting in possible preemption of critical tasks [27, 28]. I/O requests also suffer from added latency due to cross-layer traversal [29], and debugging becomes more difficult due to poor visibility across layers [30].
Furthermore, many hypervisors rely on static or semi-static resource allocation [31, 32], which cannot adapt to fluctuating workloads common in edge scenarios. This often leads to underutilization or resource contention. Additionally, inter-VM communication—even between co-located guests—typically relies on virtual network stacks, introducing unnecessary overhead [33, 34].
@@
3.3 Issues with Hypervisors in Edge Environments
Most current hypervisors adopt a monolithic architecture, which, while suitable for traditional server environments, introduces several limitations when deployed at the edge. Existing hypervisors are primarily designed for homogeneous hardware, focusing on uniform server-grade processors and devices. However, edge environments are inherently heterogeneous, featuring a diverse mix of hardware accelerators such as GPUs, FPGAs, and specialized AI chips [35, 36]. Monolithic hypervisors struggle to efficiently manage and abstract these heterogeneous resources, making it difficult to extend support for new devices and often leading to suboptimal hardware utilization [37, 35].
Moreover, the increasing complexity of hypervisor codebases has made comprehensive security verification a significant challenge. Traditional monolithic type-1 hypervisors can comprise hundreds of thousands of lines of code [38], resulting in a large attack surface that is practically infeasible to fully audit or verify, especially for safety- and security-critical applications. This challenge is particularly pronounced in edge deployments, where devices frequently operate in physically untrusted or exposed environments, amplifying the risks associated with potential vulnerabilities in the hypervisor.

### v2

# TenonOS: A Self-Generating LibOS-on-LibOS Framework for Time-Critical Embedded Operating Systems

@@ Abstract
The increasing complexity of modern embedded systems creates a fundamental conflict between the demand for sophisticated functionality and the strict constraints on resources and temporal determinism. Traditional operating system and hypervisor architectures, typically reliant on monolithic designs or rigid abstraction layers, suffer from significant resource bloat and unpredictable scheduling behaviors. These limitations render them ill-suited for time-critical applications where minimizing latency and jitter is paramount.
To address these challenges, we propose TenonOS, a demand-driven, self-generating, and lightweight operating system framework for time-critical embedded systems that fundamentally rethinks and reconstructs both the hypervisor and OS architectures. TenonOS introduces a novel LibOS-on-LibOS architecture, in which both hypervisor and OS functionalities are decomposed into fine-grained, reusable micro-libraries. Unlike static legacy systems, TenonOS employs a generative orchestration engine that dynamically composes these modules to synthesize a customized runtime environment tailored specifically to the application’s criticality level, timing constraints, and resource profile.
At the core of this framework are two synergistic components: Mortise, a minimalist micro-hypervisor, and Tenon, a real-time LibOS. Mortise provides lightweight resource isolation and eliminates the ”double scheduler” overhead common in virtualized environments, while Tenon delivers precise, deterministic task management. By generating only the necessary software stack for each workload, TenonOS eliminates redundant layers, minimizes the Trusted Computing Base (TCB), and maximizes system responsiveness. Extensive evaluations demonstrate that TenonOS achieves superior real-time performance with a 40.28% improvement in scheduling latency, maintains an ultra-compact memory footprint of 361 KiB, and exhibits high adaptability. These results validate TenonOS as an ideal foundation for next-generation, resource-constrained embedded systems requiring strict temporal guarantees.
@@
LibOS , Hypervisor , Embedded virtualization , Real-time operating systems
journal: Nuclear Physics B
@@
[1]organization=College of Software Technology, Zhejiang University, city=Ningbo, country=China
@@
[2]organization=Yingyi Technology Co., Ltd., city=Hangzhou, country=China
[…]
[3]organization=College of Computer Science and Technology, Zhejiang University, city=Hangzhou, country=China
[…]
[4]organization=State Key Laboratory of Mathematical Engineering and Advanced Computing, city=Wuxi, country=China
[…]
[5]organization=Chinese Academy of Engineering, city=Beijing, country=China
@@ 1 Introduction
Operating systems (OSs) form the backbone of the information industry, providing the essential interface between hardware and software. At the foundational level, they manage and coordinate hardware resources; at a higher level, they support the development and evolution of complex application ecosystems [peterson1985operating, tanenbaum1997operating]. The enduring success of operating systems stems from their ability to offer robust, adaptable, and scalable environments for both developers and users. However, the rapid pace of technological innovation and the explosive growth of interconnected devices are increasingly straining traditional OS architectures [armbrust2010view, satyanarayanan2001pervasive].
@@
These limitations are particularly pronounced in the context of modern embedded environments, where applications are highly dynamic, heterogeneous, and resource constrained [satyanarayanan2017emergence, shi2016edge]. The proliferation of smart devices, sensors, and interconnected systems has resulted in a fragmented and diverse network landscape [gubbi2013internet]. Unlike traditional servers, these systems, ranging from industrial robots to autonomous vehicles, operate in close proximity to physical processes, demanding strict real-time responsiveness and deterministic behavior [shi2016edge]. Operating systems in these settings must manage diverse hardware platforms (e.g., multi-core SoCs with accelerators) while ensuring strong security and isolation under stringent power and memory constraints. Consequently, conventional OS architectures, originally designed for stable, centralized computing, often fail to accommodate the fragmented and time-sensitive nature of these modern workloads [zahran2019heterogeneous, zhang2024operating].
@@
A particularly pressing challenge in these complex environments is the management of mixed-criticality workloads. Many advanced embedded applications require distinct operating environments to coexist on a single device. For instance, autonomous vehicles commonly rely on a Real-Time Operating System (RTOS) to handle safety-critical control loops, while simultaneously running a General-Purpose Operating System (GPOS) for infotainment and high-level perception [wulf2021survey, mounir2019hardware, barletta2022achieving]. This architectural complexity underscores the growing difficulty in manually integrating and optimizing OS components. The demand for efficient coordination between these diverse subsystems highlights the urgent need for a framework capable of automatically tailoring the operating system structure to meet specific hardware and application requirements.
@@
To address these needs, virtualization technologies have become a mainstream solution, enabling multiple isolated environments to run on shared hardware. By abstracting the underlying physical resources, virtualization improves flexibility and scalability, and provides strong isolation between different OS instances [shi2016edge, morabito2018consolidate]. However, virtualization is not without its drawbacks. Different layers—such as hypervisors, container runtimes, and guest operating systems—often employ independent resource management strategies, leading to inefficiencies. These inconsistencies can cause resource contention, increased latency, and degraded overall system performance, which are particularly problematic for latency-sensitive and resource-constrained applications [pan2019performance, sheikhalishahi2015autonomic, zhao2018cross, desina2023evaluating, krebs2015performance, mukherjee2015resource]. Moreover, the overhead and complexity introduced by hypervisor-based solutions may compromise real-time guarantees and make system design more challenging, especially when supporting a wide variety of hardware configurations.
@@
To tackle these challenges, we propose TenonOS, a self-generating, intelligent, and lightweight operating system framework for time-critical embedded systems. TenonOS is built on a novel LibOS-on-LibOS architecture that rethinks the conventional separation between the operating system and the hypervisor. Adopting a unified library-first design, TenonOS decomposes both layers into modular micro-libraries, with each functionality implemented as a minimal, composable library. By supporting dynamic composition at runtime, TenonOS provides a flexible and efficient execution environment that can be tailored to different criticality levels and real-time requirements. This approach eliminates much of the redundancy in traditional OS–hypervisor stacks, enabling more compact, scalable, and hardware-aware deployments on resource-constrained platforms.
@@
At the core of TenonOS is a two-layer architecture consisting of the hypervisor Mortise and the runtime Tenon. This structure is inspired by the natural growth of plants: Mortise serves as the fertile soil of the system, while Tenon acts as the seed from which multiple distinct operating systems can grow.
@@
Mortise is a lightweight hypervisor that decomposes traditional virtualization functionality into modular micro-libraries. It supports efficient hardware resource management, low-latency inter-VM communication, and flexible dynamic allocation policies. Tenon, built atop Mortise, is a real-time capable LibOS designed to meet the timing constraints of modern embedded applications. Unlike conventional LibOS designs, Tenon integrates real-time scheduling mechanisms that support both cooperative and preemptive models, enabling deterministic task execution for critical workloads. It also provides multi-process support, preserving compatibility with general-purpose applications while maintaining a compact, modular structure. Mortise can concurrently host both Tenon and general-purpose operating systems, such as Linux, on the same hardware platform. This architecture supports mixed-criticality workloads and reduces the need for separate compute nodes.
@@
A key feature of Mortise is its ability to manage the full lifecycle of Tenon instances, including on-demand instantiation, suspension, and termination of LibOS-based runtimes according to system workload and application requirements. By enabling fine-grained control over when and how operating system instances are created and destroyed, Mortise underpins TenonOS’s self-generating capability. Rather than relying on static runtime environments, TenonOS can dynamically assemble, launch, and retire execution contexts as needed, allowing the system to elastically adapt to workload variations while minimizing resource consumption.
@@
Together, Mortise and Tenon form a cohesive framework that bridges hypervisor-level resource control with application-level runtime flexibility. Mortise provides low-level isolation and scheduling, while Tenon delivers customizable execution environments optimized for responsiveness and scalability. This design aligns well with the diverse and performance-sensitive demands of resource-constrained and time-critical applications.
This paper presents the design and implementation of TenonOS, a generative operating system framework built upon a novel LibOS-on-LibOS architecture tailored for heterogeneous computing environments. Our main contributions are summarized as follows:
  • 1.
    LibOS-on-LibOS system architecture. We propose TenonOS, a unified LibOS-on-LibOS architecture that decomposes both the hypervisor and the operating system into a shared, capability-based micro-library pool. This design collapses the traditional OS–hypervisor layering and enables cross-layer code reuse and coherent resource management, providing a foundation for time-critical embedded deployments.
  • 2.
    Mortise: a minimal, dual-mode LibOS hypervisor. We design and implement Mortise, a LibOS-based type-1 hypervisor that factorizes virtualization services into micro-libraries and supports both static CPU partitioning and dynamic, hypervisor-driven scheduling modes. Mortise provides strong isolation, low-overhead inter-VM communication, and a substantially smaller trusted computing base than existing embedded hypervisors.
  • 3.
    Tenon: a real-time, multi-process LibOS for time-critical workloads. We extend a cloud-oriented LibOS stack into Tenon, a real-time capable LibOS that integrates a priority-based preemptive scheduler, fine-grained interrupt handling, and lightweight multi-process support. Our experiments show that Tenon achieves lower interrupt and scheduling latency than state-of-the-art RTOSes such as Zephyr and RT-Thread, while preserving near bare-metal behavior even when virtualized by Mortise.
  • 4.
    Open-source prototype and evaluation on ARM64 platforms. We provide an open-source implementation of TenonOS, including Tenon (https://gitee.com/tenonos/tenon.git) and Mortise (https://gitee.com/tenonos/mortise.git). Our prototype on ARM64 SoCs demonstrates a total code size of 11.3 K SLoC, a memory footprint of about 361 KiB, sub-40 ms Tenon boot time, and negligible virtualization overhead on real-time metrics even when co-located with Linux.
2 Background and Motivation
@@
The landscape of embedded systems [hamdan2020edge, gamatie2009case] has evolved dramatically, transitioning from simple control tasks to complex, data-intensive workloads. This distributed computing model brings computation and data storage closer to data sources, enabling real-time processing for applications such as autonomous vehicles [huang2023unmanned, zhou2020mobile, abrar2021energy], industrial automation [stankovski2020impact, chalapathi2021industrial, weibin2019three], and smart cities [khan2020edge, taleb2017mobile]. The unique characteristics of edge environments, including heterogeneous hardware architectures, resource-constrained devices, and geographically distributed deployments, pose significant challenges for traditional computing approaches.
@@
The evolution of operating systems has historically been shaped by the needs of centralized computing environments. Classical operating system architectures, developed during the mainframe and personal-computing eras, were designed under assumptions of relatively homogeneous hardware and abundant resources [peterson1985operating, engler1995exokernel]. These systems typically adopt monolithic designs that abstract hardware resources through rigid interfaces. While effective in desktop settings, these architectures exhibit substantial limitations in embedded scenarios. Crucially, traditional monolithic kernels lack the flexibility to efficiently isolate diverse workloads—such as mixing real-time tasks with general-purpose computing—without introducing unpredictable interference and resource contention.
Modern embedded applications demand strict real-time guarantees and high-throughput responsiveness, necessitating effective workload consolidation [liu2019edge, mach2017mobile]. To meet these requirements, virtualization and containerization are commonly adopted to provide isolation and deployment flexibility [morabito2018consolidate]. However, these traditional solutions often incur prohibitive overheads in resource-constrained embedded environments. They typically rely on deep software stacks that involve trapping and emulating instructions or maintaining duplicate kernel structures [chiueh2005survey]. This architectural redundancy consumes valuable CPU cycles and memory, directly undermining the deterministic timing behavior required by critical tasks. Consequently, existing virtualization approaches struggle to strike the right balance between isolation and efficiency, highlighting the need for a lightweight framework capable of fine-grained, hardware-aware resource management.
@@ 2.1 Issues with Monolithic Architectures
In traditional monolithic operating system and virtualization architectures, all core functionalities are tightly coupled within a large and complex codebase. Although this design was adequate in earlier eras with relatively homogeneous and stable hardware, it exhibits significant limitations in today’s heterogeneous and rapidly evolving embedded environments. Figure 1 shows that, across mainstream monolithic systems such as Linux and FreeBSD, architecture-specific code is spread over many subsystems and often reaches tens of thousands of lines per module. This broad and uneven distribution of adaptation code implies that supporting a new processor architecture or modifying an existing one requires invasive changes to multiple kernel components, increasing engineering effort and the probability of regressions. The driver ecosystem further amplifies these issues. As illustrated in Figure 2, monolithic kernels maintain a very large number of drivers, each with many architecture-specific dependencies and frequent commits over time. This tight coupling between drivers and low-level architectural interfaces makes hardware evolution costly to support and complicates long-term maintenance, since even small architectural changes can trigger cascading updates across hundreds of drivers. Consequently, monolithic architectures struggle to provide efficient, low-risk evolution in the face of diverse accelerators, rapidly changing hardware platforms, and dynamic workloads, and are increasingly inadequate for modern embedded computing environments.
Figure 1: Average architecture-specific code size per module for each operating system . Bars are grouped by module on a logarithmic scale, showing how much architecture adaptation effort each OS concentrates in different subsystems.
(a) Average number of architecture-specific dependencies per driver (left axis) and total number of drivers (right axis) for each operating system.
(b) Average total commits per driver (left axis) and average yearly commits per driver (right axis, both in log scale) across operating systems.
Figure 2: Comparison of driver ecosystem size, architectural dependencies, and maintenance activity across different operating systems.
@@ 2.2 The Case for Virtualization-OS Structures
A key limitation stems from the layered architecture of traditional systems, where both the hypervisor and the OS independently manage hardware resources. This separation leads to duplicated functionality and conflicting policies. For example, virtual memory is maintained both at the hypervisor and guest OS levels [waldspurger2002memory], and CPU scheduling occurs twice—first by the hypervisor for vCPUs, then by the OS for processes—resulting in possible preemption of critical tasks [menon2005diagnosing, govindan2007xen]. I/O requests also suffer from added latency due to cross-layer traversal [menon2006optimizing], and debugging becomes more difficult due to poor visibility across layers [cherkasova2007virtual].
Furthermore, many hypervisors rely on static or semi-static resource allocation [martins2020bao, guthrie2013vmware], which cannot adapt to fluctuating workloads common in embedded scenarios. This often leads to underutilization or resource contention. Additionally, inter-VM communication—even between co-located guests—typically relies on virtual network stacks, introducing unnecessary overhead [barker2012empirical, ivanovic2018performance].
@@
2.3 Issues with Hypervisors in Embedded Environments
Most current hypervisors adopt a monolithic architecture, which, while suitable for traditional server environments, introduces several limitations when deployed in embedded environments. Existing hypervisors are primarily designed for homogeneous hardware, focusing on uniform server-grade processors and devices. However, embedded environments are inherently heterogeneous, featuring a diverse mix of hardware accelerators such as GPUs, FPGAs, and specialized AI chips [ottaviano2024omnivisor, zha2021hetero]. Monolithic hypervisors struggle to efficiently manage and abstract these heterogeneous resources, making it difficult to extend support for new devices and often leading to suboptimal hardware utilization [gupta2015heterovisor, ottaviano2024omnivisor].
Moreover, the increasing complexity of hypervisor codebases has made comprehensive security verification a significant challenge. Traditional monolithic type-1 hypervisors can comprise hundreds of thousands of lines of code [steinberg2010nova], resulting in a large attack surface that is practically infeasible to fully audit or verify, especially for safety- and security-critical applications. This challenge is particularly pronounced in embedded environments, where devices frequently operate in physically untrusted or exposed environments, amplifying the risks associated with potential vulnerabilities in the hypervisor.